# Grid-based Multi-Robot Coverage Path Planning

### v1

@@ Abstract
Coverage Path Planning (CPP) addresses the challenge of planning paths to completely cover a designated workspace and is crucial for various robotic applications, including indoor cleaning, inspection, and outdoor tasks such as automated harvesting and environmental monitoring. Multi-Robot Coverage Path Planning (MCPP) extends this by coordinating multiple robots to enhance coverage efficiency and system robustness. This paper focuses on the grid-based MCPP problem, where the workspace is represented as a four-way connected 2D grid graph. Unlike traditional approaches that rely on the k-Traveling Salesman Problem (k-TSP) or Spanning Tree Coverage (STC), we reformulate MCPP directly on the edge-weighted graph to be covered and incorporate inter-robot conflicts and turning costs, reflecting more practical considerations in real-world applications. We propose several key improvements over our previous work: a reformulated Extended-STC (ESTC) with local optimizations, a refined local search framework (LS-MCPP), and an adaptive multi-goal planning approach combined with Priority-Based Search (PBS) for conflict resolution. Our proposed pipeline scales efficiently to large grid graphs (up to $256\times 256$) and up to 100 robots, demonstrating superior solution quality and success rates within about 18 minutes. We also validate our approach through a multi-robot system deployment, showcasing its practical effectiveness in real-world scenarios.
@@
Multi-Robot System, Automated Planning.
@@ I Introduction
Coverage Path Planning (CPP) addresses the problem of determining a path that fully covers a designated workspace [galceran2013survey]. This problem has been extensively studied for a variety of robotic applications, ranging from indoor tasks like vacuum cleaning [wang2024apf] and inspection [almadhoun2018coverage], to outdoor activities such as automated harvesting [chen2024optimizing], planetary exploration [santra2024risk], and environmental monitoring [sudha2024coverage]. Multi-Robot Coverage Path Planning (MCPP) extends traditional CPP by coordinating the paths of multiple robots to collectively cover a workspace, thereby enhancing both the efficiency of the coverage task [tang2022learning] and the robustness of the overall system [sun2021ft]. Developing efficient and robust MCPP algorithms is crucial for the widespread deployment of multi-robot systems across various real-world applications. In this paper, we focus on a specific MCPP formulation known as grid-based MCPP, where the workspace is represented as a four-way connected 2D grid graph. The objective is for a team of robots to jointly cover all vertices of the graph, a problem analogous to the NP-hard k-Traveling Salesman Problem (k-TSP)[francca1995m]. However, solving k-TSP in the context of grid-based MCPP is challenging due to the large number of robots and the extensive graph size, which often comprises tens of thousands of vertices. Consequently, existing grid-based MCPP methods leverage the grid structure to find solutions that are constrained to the four-way connected grid space. Specifically, Spanning Tree Coverage (STC)[gabriely2001spanning] is used to generate circumnavigating coverage paths for individual robots, which sequentially follow one side of the spanning edges on a hypergraph (also known as the terrain graph) of the decomposed grid. The MCPP is then solved either by segmenting the STC path and distributing the segments among the robots[hazon2005redundancy, tang2021mstc], or by identifying the covering subtrees and generating separate STC paths for each robot [zheng2007robot, zheng2010multirobot]. For simplicity, the prefix ”grid-based” is omitted throughout the paper when referring to MCPP unless specified otherwise.
Figure 1: Simulation of 100 robots cooperatively performing a coverage task on the $256\times 256$ grid map NewYork1 [sturtevant2012benchmarks] without conflicts using our proposed planning pipeline, which reduces task completion time (i.e., makespan) by 42% compared to the baseline and resolves 556 conflicts within 18 minutes. Left: Task progress at one-third completion. Right: Task is completed.
We reformulate MCPP directly on the edge-weighted graph to be covered, explicitly considering inter-robot conflicts (i.e., collisions) to enable broader applicability in real-world robotics scenarios. The edge weights can represent various meaningful metrics, such as traversal costs on uneven terrains [qiu2024terrain] or the importance of task-specific edges that may require additional traversal time [freda20193d]. Additionally, we account for turning costs when evaluating path costs on the graph —– a critical factor in real-world coverage applications [vandermeulen2019turn, lu2022tmstc, ramesh2022optimal], as it demands additional execution time for robots. This problem setup is significantly more challenging than in the existing literature, where traversal costs are defined on hypergraph edges without considering turning costs and inter-robot conflicts. To address these challenges, we introduce several key improvements over our previous work [tang2024large], which initially computes a set of covering subgraphs for the target graph and subsequently generates a single-robot circumnavigating path for each subgraph, ensuring that each robot efficiently covers its assigned region. First, we describe the multi-robot system and formulate the problem in Sec.III. Next, we enhance the Extended-STC (ESTC) method by encoding cost differences between circumnavigating paths based on whether each hypergraph edge is used, and introduce two local optimizations (Sec.IV). With ESTC, we enhance our local search framework (LS-MCPP) to iteratively improve an initial solution. This refinement incorporates additional vertex-wise operators that facilitate better exploration of the solution space while initially ignoring inter-robot conflicts. (Sec.V). Finally, we propose an adaptive approach as a low-level multi-goal planner for each robot, combined with Priority-Based Search (PBS)[ma2019searching] as a high-level planner to resolve pairwise conflicts, together efficiently deconflicting the LS-MCPP solution (Sec.VI). Our pipeline efficiently scales to instances involving up to $256\times 256$ grid graphs and 100 robots, achieving superior solution quality and success rates compared to existing approaches within a runtime of approximately 15 minutes (Sec.VII). Moreover, we deploy a multi-robot system to demonstrate the effectiveness of our pipeline in real-world robotics applications (Sec. VII-E).
@@ II-A Multi-Robot Coverage Path Planning (MCPP)
MCPP focuses on finding a set of paths for a team of robots to achieve complete coverage of a given workspace. Following the taxonomy of single-robot CPP methods [Tomaszewski-2020-125840], we review three types of MCPP methods according to their levels of decomposition on the target workspace.
Grid-Based Methods: Grid-based methods represent the workspace as a grid graph [kapoutsis2017darp, azpurua2018multi, li2023sp2e], allowing the application of various graph algorithms. One common approach involves decomposing the graph for the robots [zheng2007robot, zheng2010multirobot, tang2023mixed] and using the Spanning Tree Coverage (STC) algorithm [gabriely2001spanning] to generate single-robot coverage paths. As mentioned in Sec. I, these methods operate on the hypergraph of the workspace, where each hypergraph vertex is decomposed into four smaller vertices. Rather than segmenting the graph directly, they decompose the hypergraph because STC relies on a spanning tree of the hypergraph to generate the single-robot coverage paths. Our previous work [tang2024large] proposed an alternative approach that operates directly on the graph to be covered, utilizing an extended version of STC to generate single-robot coverage paths on any subgraph. Unlike the hierarchical approach mentioned above, another type of approach [hazon2005redundancy, tang2021mstc] first generates a single-robot coverage path and then divides it into multiple segments, which are subsequently distributed among the robots. A significant assumption in these grid-based methods is that the hypergraph is vertex-weighted, with the weight of each hypergraph vertex evenly distributed across the four decomposed vertices. The cost of each coverage path is evaluated by summing the weights of all the decomposed vertices along the path, which may not accurately reflect the actual traversal costs for the robots, as these weights heavily constrain them. A more general and effective practice is to directly define the traversal cost based on the weights of the edges traversed by the robots, as in our formulation in our formulation in Sec. III-A), which, in the case of grid-based MCPP, corresponds to the edges of the graph to be covered.
@@
Cellular Decomposition Methods: Cellular decomposition methods [rekleitis2008efficient, xu2014efficient, karapetyan2017efficient] divide the workspace into sub-regions by detecting geometric critical points, using techniques such as trapezoidal decomposition [latombe1991exact], boustrophedon decomposition [choset2000coverage], and Morse decomposition [acar2002morse]. After the division, a back-and-forth patterned path is generated to cover each sub-region, with possible optimization of the path alignment orientation [bochkarev2016minimizing]. While efficient, these methods are less suited for obstacle-rich or non-rectil inear workspaces due to their reliance on geometric partitioning. Additionally, they assume uniform traversal costs across the workspace, making them inapplicable in scenarios where robots require nonuniform time to move or perform specific coverage tasks.
@@
Global Methods: Global methods are primarily studied and applied in additive manufacturing [yang2002equidistant, ren2009combined, gibson2021additive] (also known as layered fabrication), where a 3D object is printed layer by layer. Each layer represents a coverage problem, requiring the planning of a ”space-filling” curve for the nozzle to cover the 2D plane. Due to the path curvature requirements in these applications, global methods do not explicitly decompose the workspace; instead, they directly generate paths to cover the workspace, minimizing path curvature and discontinuity. Recent works have extended the connected Fermat spiral [zhao2016connected] to reduce energy consumption for CPP on undulating terrains [wu2019energy] and to minimize the makespan for general MCPP tasks [tang2024multi]. The coverage quality (e.g., coverage and overlapping ratios) of global methods depends on the nature of the generated space-filling curve, often requiring additional trajectory optimization as a post-processing step. A further challenge for global MCPP methods is task distribution among robots for makespan minimization, as the space-filling curve typically forms a closed loop with two endpoints.
@@ II-B Multi-agent Path Finding (MAPF)
MAPF seeks to generate a set of plans for multiple agents to move from start to goal locations without conflicts (i.e., collisions) with other agents. For detailed surveys on MAPF, we refer readers to [stern2019multi, stern2019multi2, ma2022graph]. Here, we review the multi-goal variants of MAPF most relevant to this paper.
A recent study formulates an ordered multi-goal MAPF problem (OMG-MAPF) [mouratidis2024fools] to find plans for multiple agents so that each agent visits an ordered set of goal locations without conflicts. The order of goal locations for each agent is known as the total order. A related problem, MAPF-PC [zhang2022multi], considers precedence constraints on pairs of goal locations for each agent, meaning one goal location must be visited before the other. This precedence constraint is referred to as the partial order, and partial orders may only apply to specific pairs of agents in MAPF-PC. A set of partial orders can be converted into a total order via topological sorting, and a total order can also be extracted from a set of partial orders. For both OMG-MAPF and MAPF-PC, it is assumed that all actions have unit costs and that agents operate on a clock discretized into unit time steps. Thus, these methods are not directly applicable to graphs with nonuniform edge weights.
Another line of work on multi-goal variants, called MG-MAPF [surynek2021multi, ren2021ms, ren2022conflict, tang2024mgcbs], considers an unordered set of goal locations for each agent, with the additional objective of determining the optimal global ordering of goal locations (i.e., optimal goal sequencing) to minimize an objective function. Although one might formulate MCPP (as defined in Problem III) as an MG-MAPF problem, co-optimizing the goal sequencing for the vertices along each coverage path might not yield a better sequence than the original coverage path within a limited planning time, given the combinatorial complexity of co-optimization. Intuitively, as the robot must start and return to its root vertex in MCPP, most conflicts can be resolved by short-duration wait actions without the need to re-sequence the vertices along a coverage path.
The Multi-agent Pickup and Delivery (MAPD) problem [ma2017lifelong, ma2019lifelong, li2021lifelong, xu2022multi] can also be considered a multi-goal MAPF variant, where a system continuously assigns new tasks to agents. However, agents are unaware of each task’s pickup and delivery locations before the task is assigned, making it more difficult to address the goal sequencing problem.
III Problem Formulation
In this section, we first describe the multi-robot system with several key assumptions. Upon this, we address the MCPP problem through a progressive approach by solving a relaxed version of MCPP that disregards inter-robot conflicts and then resolving the conflicts among the generated coverage paths.
III-A Description of the Multi-Robot System

### v2

# Large-Scale Multi-Robot Coverage Path Planning on Grids with Path Deconfliction

@@ Abstract
We study Multi-Robot Coverage Path Planning (MCPP) on a 4-neighbor 2D grid $G$, which aims to compute paths for multiple robots to cover all cells of $G$. Traditional approaches are limited as they first compute coverage trees on a quadrant coarsened grid $\mathcal{H}$ and then employ the Spanning Tree Coverage (STC) paradigm to generate paths on $G$, making them inapplicable to grids with partially obstructed $2\times 2$ blocks. To address this limitation, we reformulate the problem directly on $G$, revolutionizing grid-based MCPP solving and establishing new NP-hardness results. We introduce Extended-STC (ESTC), a novel paradigm that extends STC to ensure complete coverage with bounded suboptimality, even when $\mathcal{H}$ includes partially obstructed blocks. Furthermore, we present LS-MCPP, a new algorithmic framework that integrates ESTC with three novel types of neighborhood operators within a local search strategy to optimize coverage paths directly on $G$. Unlike prior grid-based MCPP work, our approach also incorporates a versatile post-processing procedure that applies Multi-Agent Path Finding (MAPF) techniques to MCPP for the first time, enabling a fusion of these two important fields in multi-robot coordination. This procedure effectively resolves inter-robot conflicts and accommodates turning costs by solving a MAPF variant, making our MCPP solutions more practical for real-world applications. Extensive experiments demonstrate that our approach significantly improves solution quality and efficiency, managing up to 100 robots on grids as large as $256\times 256$ within minutes of runtime. Validation with physical robots confirms the feasibility of our solutions under real-world conditions. A project page with code, demo videos, and additional resources is available at: https://sites.google.com/view/lsmcpp.
@@
Multi-Robot System, Automated Planning, Heuristic Search.
[…]
@@ I Introduction
Coverage Path Planning (CPP) addresses the problem of determining a path that fully covers a designated workspace [1]. This problem is essential for a broad spectrum of robotic applications, from indoor tasks like vacuum cleaning [2] and inspection [3] to outdoor activities such as automated harvesting [4], planetary exploration [5], and environmental monitoring [6]. Multi-Robot Coverage Path Planning (MCPP) is an extension of CPP tailored for multi-robot systems, aiming to coordinate the paths of multiple robots to collectively cover the given workspace, thereby enhancing both task efficiency [7] and system robustness [8]. A fundamental challenge of MCPP is to balance the costs across multiple robots, often quantified by the makespan—the maximum path cost among all robots. This challenge intensifies in large-scale applications where the number of robots and the size of the workspace significantly increase.
Figure 1: Grid-based CPP and MCPP: black circles, dark gray squares, solid lines, and dashed lines represent vertices of $G$, vertices of $\mathcal{H}$, coverage paths on $G$, and spanning edges of $\mathcal{H}$, respectively. Starting at $r$, the robot(s) need to collectively cover each vertex of $G$ and return to $r$, with a cost of $1$ to traverse between adjacent vertices of $G$. (a) Single-robot coverage path generated by STC. (b) Suboptimal $2$-robot coverage paths with a makespan of $8$, resulting from 2 coverage trees that minimize the maximum tree weights. (c) Optimal $2$-robot coverage paths with a makespan of $6$.
This article tackles MCPP on a 4-neighbor 2D (edge-)weighted grid graph $G$. The problem is to compute paths for a given set of $k$ robots, each starting from and returning to their given root vertices, to jointly cover all vertices of $G$, a problem analogous to the NP-hard Multiple Traveling Salesman Problem (m-TSP) [9] and Vehicle Routing Problem (VRP) [10]. Existing grid-based MCPP methods [11, 12, 13, 14, 15] operate on a coarsened grid $\mathcal{H}$, derived from $G$ by grouping every $2\times 2$ adjacent vertices, to compute coverage trees that jointly cover all vertices of $\mathcal{H}$. These methods then apply the Spanning Tree Coverage (STC) [16] paradigm to generate coverage paths on $G$ by circumnavigating the computed coverage trees. However, operating exclusively on the coarsened grid $\mathcal{H}$ has two significant disadvantages. First, it fails in cases where $\mathcal{H}$ is incomplete—that is, when any $2\times 2$ blocks contain obstructed grid cells absent from $G$ [^1] The grid graph $G$ is technically a vertex-induced subgraph of an infinite grid. For clarity and consistency, we refer to the vertices in the infinite grid as ”grid cells” throughout this paper. These grid cells can be obstructed, resulting in their absence as vertices in $G$. This distinction helps articulate the difference between the physical presence of cells in the infinite grid and their representation (or lack thereof) in the subgraph $G$. Such terminology is particularly useful in scenarios where contracting a $2\times 2$ block of grid cells does not yield a complete set of four vertices, or a “hypervertex”, due to the obstruction or absence of one or more grid cells. This approach enables us to discuss the structure and topology of $G$ more precisely, especially in the context of MCPP problems where the presence or absence of grid cells affects path planning and coverage strategies. Second, even optimal coverage trees on $\mathcal{H}$ do not necessarily result in an optimal MCPP solution (as illustrated in Fig. (1b) and (1c)), as evidenced by an asymptotic suboptimality ratio of four for makespan minimization [13], since the paths derived from circumnavigating coverage trees of $\mathcal{H}$ constitute only a subset of all possible sets of coverage paths on $G$. Furthermore, existing grid-based MCPP methods neglect to resolve inter-robot conflicts [17] in their computed solutions, leading to potential collisions during execution and thus limiting their practical application in diverse real-world multi-robot scenarios.
I-A Contributions
We revolutionize solving MCPP on grid graphs, overcoming the above limitations through a two-phase approach that first systematically searches for good coverage paths directly on $G$ and subsequently resolves inter-robot conflicts in a post-processing procedure, which is illustrated in Fig. (2). Our algorithmic contribution are detailed as follows:
  1. 1.
    We propose a novel standalone paradigm called Extended-STC (ESTC), which extends STC to address coverage path-planning problems on any grid graph $G$, applicable both when the quadrant coarsening $\mathcal{H}$ is complete and incomplete. Importantly, we demonstrate that ESTC guarantees complete coverage for both single- and multi-robot settings with bounded suboptimality, rendering it an efficient and versatile solution for coverage path planning.
  2. 2.
    We design three types of specialized neighborhood operators to facilitate an effective local search process by identifying cost-efficient subgraphs of $G$ that are then used to generate coverage paths for the robots. The strategic integration of these operators significantly improves the efficiency of exploring the solution space. We then demonstrate how to combine these neighborhood operators with iterative calls to the ESTC paradigm to establish our LS-MCPP framework for solving MCPP.
  3. 3.
    We address inter-robot conflicts in the coverage paths computed by LS-MCPP or other MCPP methods by formulating the problem as a variant of Multi-Agent Path Finding [17]. This variant asks to compute conflict-free continuous-time trajectories for robots as each of them visits an ordered sequence of vertices specified by the MCPP solution. We develop new approaches for solving this variant based on existing MAPF techniques. These approaches integrate the novel Multi-Label Safe Interval Path Planning search that plans a continuous-time trajectory for an individual robot through its ordered sequence of vertices and, as a side benefit, allows our post-processing step to accommodate important kinematic constraints, such as turning costs—–a critical factor in real-world coverage applications [18, 19, 20],
We showcase the benefits of our algorithmic pipeline through extensive quantitative experiments on large-scale instances involving up to 100 robots on a $256\times 256$ grid. Key findings include:
  1. 1.
    Our ESTC paradigm results in coverage paths with significantly smaller makespans than existing STC-based paradigms and is effective on any grid graphs $G$, including those with incomplete quadrant coarsening $\mathcal{H}$, for both CPP and MCPP tasks.
  2. 2.
    Our LS-MCPP framework achieves significantly smaller makespans than state-of-the-art MCPP methods that rely on suboptimal tree coverage computations on $\mathcal{H}$ and requires far less runtime to achieve makespans comparable to or better than those achieved by the optimal tree coverage computation.
  3. 3.
    Our post-processing procedure can be applied to any MCPP solutions, including not only those generated by LS-MCPP but also those generated by existing MCPP methods, to effectively resolve conflicts between robots and accounts for turning costs, further enhancing the practicability of the solutions.
Additionally, we validate the feasibility of our approach with physical robot deployments, confirming its practical applicability in real-world robotics applications.
Figure 2: The proposed two-phase approach for MCPP in a grid graph with $k$ robots. (a) The LS-MCPP phase: Starting with an initial solution $\Pi_{0}$, the algorithm iteratively samples operators to update the subgraph set $\mathcal{G}$ and generates $k$ coverage paths using ESTC, ultimately producing a makespan-minimized solution $\Pi^{*}$. (b) The deconflicting phase: Given $\Pi^{*}$ (or any other MCPP solution) as input, Priority-Based Search (PBS) explores a tree of priority ordering $\boldsymbol{\prec}$ while invoking an adaptive low-level planner to generate conflict-free paths, considering only high-priority robot trajectories for each robot. The process terminates with a set $\mathcal{T}$ of conflict-free coverage paths.
This article significantly extends its preliminary conference version [21] by:
  1. 1.
    Providing a reformulation of MCPP directly on an (edge-)weighted grid graph $G$ and a new NP-hardness result.
  2. 2.
    Improving the original ESTC paradigm by encoding cost differences between circumnavigating paths, with a new theoretical analysis of the properties of ESTC.
  3. 3.
    Incorporating new vertex-wise operators into the original LS-MCPP framework to facilitate a more effective exploration of the solution space.
  4. 4.
    Introducing the new post-processing procedure to resolve conflicts between robots.
  5. 5.
    Expanding the numerical results and including physical robot experiments to demonstrate the practical applicability of our approach.
I-B Organization
We organize the remainder of this article as follows: We review related problems and methods in Sec. II. We present the problem formulations and complexity results in Sec. III. We introduce the ESTC paradigm in Sec. IV. We detail the specialized neighborhood operators and the LS-MCPP framework in Sec. V). We present our proposed post-processing procedure for resolving inter-robot conflicts in Sec.VI). We present our numerical results in Sec. VII and discuss our physical robot experiments in Sec. VII-D. We conclude our findings and outline future directions in Sec.VIII.
@@ II-A Multi-Robot Coverage Path Planning (MCPP)
Following the existing taxonomy on CPP [22], we categorize MCPP methods into grid-based, cellular decomposition, and global methods based on their approach to workspace decomposition.
Grid-Based Methods: Grid-based methods conceptualize the workspace as a grid graph $G$ [23, 24, 25], allowing for the application of various graph algorithms. These methods traditionally operate exclusively on a quadrant coarsened $\mathcal{H}$ derived from $G$ and fail when $\mathcal{H}$ is incomplete. They generalize the Spanning Tree Coverage (STC) paradigm for single-robot CPP in two ways. Single-tree methods[11, 14] compute a spanning tree on $\mathcal{H}$, generate a coverage path on $G$ that circumnavigate this tree, and then segment this path to distribute among multiple robots. These methods do not offer guarantees on solution quality due to their simplistic approach to path distribution. Multi-tree methods [12, 13, 15] compute multiple coverage trees, each rooted at the root vertex of a specific robot, to jointly cover all vertices of $\mathcal{H}$, and then generate the coverage paths on $G$ for the robots by circumnavigating their respective trees. In essence, multi-tree methods reduce MCPP to the NP-hard Min-Max Rooted Tree Cover problem [26, 27], yielding an asymptotic suboptimality ratio of four in makespan minimization on grid graphs. It is important to note that, unlike our direct formulation of MCPP on an edge-weighted $G$, existing grid-based methods typically formulate MCPP on a vertex-weighted $\mathcal{H}$, making a restricted assumption that the weight of each vertex in $\mathcal{H}$ is evenly distributed among its corresponding four vertices in $G$.
@@
Cellular Decomposition Methods: Cellular decomposition methods [28, 29, 30] divide the workspace into sub-regions by detecting geometric critical points, using techniques such as trapezoidal decomposition [31], boustrophedon decomposition [32], and Morse decomposition [33]. After dividing the workspace, these methods generate a back-and-forth patterned path to cover each sub-region, often optimizing the path alignment orientation to enhance coverage efficiency [34]. While these methods are effective in structured environments, their performance diminishes in areas dense with obstacles or in non-rectilinear workspaces due to their reliance on geometric partitioning. Furthermore, they typically assume uniform traversal costs across the workspace, limiting their applicability in scenarios where robots have varying movement costs or task-specific coverage demands.
@@
Global Methods: Global methods find significant application in additive manufacturing [35, 36, 37] (also known as layered fabrication), where a 3D object is printed layer by layer. Each layer poses a unique coverage problem, necessitating the planning of a space-filling curve to ensure the nozzle covers the entire 2D plane effectively. These methods focus on minimizing path curvature and discontinuities without explicitly decomposing the workspace. Recent advancements include the adaptation of the connected Fermat spiral [38] to reduce energy consumption in CPP tasks over undulating terrains [39] and to minimize the makespan in general MCPP scenarios [40]. The coverage quality of global methods largely depends on the characteristics of the space-filling curve generated, often necessitating additional trajectory optimization in a post-processing step.
@@ II-B Multi-Agent Path Finding (MAPF)
MAPF aims to compute conflict-free plans for multiple agents to move from their start vertices to their goal vertices. It is NP-hard to solve optimally for makespan minimization on grid graphs [41]. We highlight only relevant MAPF variants where each agent is given multiple goal vertices and direct readers interested in broader MAPF discussions to detailed surveys [17, 42]. OMG-MAPF [43] specifies a total order on the goal vertices given to each agent, while MAPF-PC [44] specifies precedence constraints—some goal vertex must be visited before another. MG-MAPF [45, 46, 47] does not specify a specific visiting order for the goal vertices given to each agent but instead aims to solve the goal sequencing problem jointly with MAPF. Multi-agent Pickup and Delivery (MAPD) [48, 49, 50] does not pre-assign goal vertices to agents and aims to solve the goal allocation and sequencing problem jointly with MAPF. All the above MAPF variants assume unit action costs and discrete time steps and do not directly apply to the problem of deconflicting MCPP solutions on graphs with nonuniform edge weights.
III MCPP and Its NP-Hardness